OpenStudy (anonymous):

A bag contains seven marbles, three red and four green. If three marbles are drawn from the bag at random, what is the probability that all three will be red? (a) 3/35 (b) 3/7 (c) 1/35 (d) 3/4 (e) 4/7

OpenStudy (anonymous):

3/7

OpenStudy (anonymous):

all three means first red, second red, third red

OpenStudy (anonymous):

no it is not \(\frac{3}{7}\)

OpenStudy (anonymous):

Oh yea u have to add them whoops

OpenStudy (anonymous):

probability first one is red is \(\frac{3}{7}\) then the probability the second one is red given that the first one is red is \(\frac{2}{6}\) because there are 2 red ones left, and 6 in the bag finally the probability that the third is red given the first two are is \(\frac{1}{5}\) because there are now 5 left in the bag and 1 red

OpenStudy (anonymous):

Probability of getting the first is 3/7

OpenStudy (anonymous):

My mistake didn't fully read question

OpenStudy (anonymous):

you get your answer by multiplying these together i.e. \[\frac{3}{7}\times \frac{2}{6}\times \frac{1}{5}\]

OpenStudy (anonymous):

1/35

OpenStudy (anonymous):

that is what i get, yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Probability of getting the first is 3/7

OpenStudy (anonymous):

yw